The source of bitterness in dark chocolate is the compound theobromine, an alkaloid present in cocoa beans. In a sample of dark chocolate there are 6.42×10^20 molecules of theobromine present.

How many milligrams of theobromine are present in the sample? The molecular formula for theobromine is C7H8N4O2

.


1
Expert's answer
2022-02-24T01:36:03-0500
"n=\frac{N}{N_A}= \frac{6.52\times10^{20}}{6.022\times 10^{23}}=1.08\times10^{-3}mol""m=M\times n"

"M(C_7H_8N_4O2)=180.164 \frac{g}{mol}"




"m(C_7H_8N_4O_2)= 180.164\times1.08\times10^{-3}= 0.195 g =195 mg"